OpenStudy (anonymous):

Problem for you. If [x] represents the greatest integer less than or equal to x. Find the number of integral solutions of [x/101] = [x/99]

OpenStudy (experimentx):

[x] is floor function right??

OpenStudy (anonymous):

Umm. Yes. Is[1.4]= 1 ?

OpenStudy (experimentx):

x=0-98 x=101-(99x2-1) x=101x2-(99x3-1) x=(101x3) - (99x4-1) ..

OpenStudy (experimentx):

that would be positive integer solutions ... I think we need for negative side too.

OpenStudy (anonymous):

I'm not getting what youre doing. A little explanation?

OpenStudy (experimentx):

I'm finding the domain of x. first case, 0 to 98, then 101 to (99x2-1) ... and so on. The domain is shrinking.

OpenStudy (experimentx):

is answer 2x2450?

OpenStudy (anonymous):

Close. Very close.

OpenStudy (experimentx):

4902 ?? i guess 4902 because it is symmetric both on negative and positive.

OpenStudy (anonymous):

Nope:)

OpenStudy (experimentx):

is it 4901??

OpenStudy (experimentx):

possibly 4903 adding the last number on both sides plus adding 0

OpenStudy (anonymous):

HAha. It's 5000.

OpenStudy (experimentx):

I see http://www.wolframalpha.com/input/?i=%282%2B4%2B..%2B100%29+%2B+%282%2B4%2B6%2B..%2B98%29 on the negative side, [-0.5] = 0 .. that gives extra hundred on the negative side.

OpenStudy (anonymous):

Excellent. Well done.As usua. :)